How thick of a concrete wall is required to stop a 6lb round fired at mach six?


1
Expert's answer
2020-10-19T13:21:55-0400

solution

mass of round(m)=6 lb=2.72 kg

velocity(u)=7350.26 km/h=2041.7 ms

by first equation of motion

v=u - at

v=0 m/s

u=at

and

"v^2=u^2-2as"


"s=\frac{u^2}{2a}"


"s=\frac{ut}{2}"


"s=1020.85t\space m"


so thickness required is 1020.85t meter.